# Towards Principled Multi-Agent Task Agnostic Exploration

### v1

@@ 5 Trust Region for Exploration in Practice
MixtureJointDisjointUniform
(a) Joint Entropy.
(b) Mixture Entropy.
Figure 2: Single-trial Joint and Mixture Entropy induced by mixture, joint or disjoint objective optimization along a $T=50$ horizon. (Right) State Distributions of two agents induced by different learned policies. We report the average and 95% c.i. over 4 runs.
(c) Mixture
(d) Joint
(e) Disjoint
(f) Uniform
The main idea then follows from noticing that the surrogate function in Eq. (5.1) consists of an Importance Sampling (IS) estimator (Owen, 2013), and it is then possible to optimize it in a fully decentralized and off-policy manner, similarly to what was done in Metelli et al. (2020) for MDPs and in Mutti & Restelli (2020) for convex MDPs. More specifically, given a pre-specified objective of interest $\zeta_{1}\in\{\zeta_{1}^{\infty},\zeta_{1}^{i},\tilde{\zeta}_{1}\}$, agents sample $N$ trajectories $\{(\mathbf{s}_{n},\mathbf{a}_{n})\}_{n\in[N]}$ from the environment by following a (joint) policy with parameters $\bm{\theta}_{0}=(\theta^{i}_{0},\theta^{-i}_{0})$. They then compute the values of the objective for each trajectory, building separate datasets $\mathcal{D}^{i}=\{(\mathbf{s}^{i}_{n},\mathbf{a}^{i}_{n}),\zeta_{1}^{n}\}_{n\in[N]}$. Each agent uses her dataset to compute the Monte-Carlo approximation of the Surrogate Function, namely:
^$\vspace{-0.2cm}\hat{\mathcal{L}}^{i}(\theta^{i}_{h}/\theta^{i}_{0})=\frac{1}{N}\sum_{n\in[N]}\rho^{i,n}_{\theta^{i}_{h}/\theta^{i}_{0}}\zeta^{n}_{1},\vspace{-0.01cm}$
where $\rho^{i,n}_{\theta^{i}_{h}/\theta^{i}_{0}}=\prod_{t\in[T]}\pi^{i}_{\theta^{i}_{h}}(\mathbf{a}^{i}_{n}[t]|\mathbf{s}^{i}_{n}[t])/\pi^{i}_{\theta^{i}_{0}}(\mathbf{a}^{i}_{n}[t]|\mathbf{s}^{i}_{n}[t])$ and $\zeta^{n}_{1}$ is the plug-in estimator of the entropy based on the empirical measure $d_{1}$ (Paninski, 2003). Finally, at each off-policy iteration $h$, each agent updates its parameter via gradient ascent $\theta^{i}_{h+1}\leftarrow\theta^{i}_{h}+\eta\nabla_{\theta^{i}_{h}}\hat{\mathcal{L}}^{i}(\theta^{i}_{h}/\theta^{i}_{0})$ until the trust-region boundary is reached, i.e., when it holds $D_{\mathrm{KL}}(\pi^{i}_{\tilde{\theta}^{i}}\|\pi^{i}_{\theta^{i}})>\delta.$ The psudo-code of TRPE is reported in Algorithm 1.
Remark.
One should note that TRPE is a multi-agent decentralized algorithm and it explicitly addresses task-agnostic exploration objectives, however the algorithmic blueprint is of independent interest since it is both able to address any convex functional $\mathcal{F}(\cdot)$, and it is valid in the single agent case as well.
6 Proof of Concept Experiments
In this section, we provide some empirical validations of the findings discussed so far. Especially, we aim to answer the following questions: (a) Is Algorithm 1 actually capable of optimizing finite-trials objectives? (b) Do different objectives enforce different behaviors, as expected from Section 3? (c) Does the clustering behavior of mixture objectives play a crucial role? If yes, when and why?
Throughout the experiments, we will compare the result of optimizing finite-trial objectives, either joint, disjoint, mixture ones, through Algorithm 1 via fully decentralized policies. The experiments will be performed with different values of the exploration horizon $T$, so as to test their capabilities in different exploration efficiency regimes. [^6] The exploration horizon $T$, rather than being a given trajectory length, has to be seen as a parameter of the exploration phase which allows to tradeoff exploration quality with exploration efficiency. The full implementation details are reported in Appendix B.
Experimental Domains.
The experiments were performed on two domains. The first is a notoriously difficult multi-agent exploration task called secret room (MPE, Liu et al., 2021b), [^7] We highlight that all previous efforts in this task employed centralized policies. We are interested on the role of the entropic feedback in fostering coordination rather than full-state conditioning, then maintaining fully decentralized policies instead. referred to as Env. (i). In such task, two agents are required to reach a target while navigating over two rooms divided by a door. In order to keep the door open, at least one agent have to remain on a switch. Two switches are located at the corners of the two rooms. The hardness of the task then comes from the need of coordinated exploration, where one agent allows for the exploration of the other. The second is a simpler exploration task yet over a high dimensional state-space, namely a 2-agent instantiation of Reacher (MaMuJoCo, Peng et al., 2021), referred to as Env. (ii). Each agent corresponds to one joint and equipped with decentralized policies conditioned on her own states. In order to allow for the use of plug-in estimator of the entropy (Paninski, 2003), each state dimension was discretized over 10 bins.
MixtureJointDisjointRandom Initialization
(a) MA-TRPO with TRPE Pre-Training (Env. (i), $T=150$).
(b) MA-TRPO with TRPE Pre-Training (Env. (i), $T=50$).
(c) MA-TRPO with TRPE Pre-Training (Env. (ii), $T=100$).
Figure 3: Effect of pre-training in sparse-reward settings.(left) Policies initialized with either Uniform or TRPE pre-trained policies over 4 runs over a worst-case goal. (rigth) Policies initialized with either Zero-Mean or TRPE pre-trained policies over 4 runs over 3 possible goal state. We report the average and 95% c.i.
Task-Agnostic Exploration.
Algorithm 1 was first tested in her ability to address task-agnostic exploration per se. This was done by considering the well-know hard-exploration task of Env. (i). The results are reported in Figure 2(b) for a short exploration horizon $(T=50)$. Interestingly, at this efficiency regime, when looking at the joint entropy in Figure 2(a), joint and disjoint objectives perform rather well compared to mixture ones in terms of induced joint entropy, while they fail to address mixture entropy explicitly, as seen in Figure 2(b). On the other hand mixture-based objectives result in optimizing both mixture and joint entropy effectively, as one would expect by the bounds in Th. 4.1. By looking at the actual state visitation induced by the trained policies, the difference between the objectives is apparent. While optimizing joint objectives, agents exploit the high-dimensionality of the joint space to induce highly entropic distributions even without exploring the space uniformly via coordination (Fig. 2(d)); the same outcome happens in disjoint objectives, with which agents focus on over-optimizing over a restricted space loosing any incentive for coordinated exploration (Fig.2(e)). On the other hand, mixture objectives enforce a clustering behavior (Fig.2(e)) and result in a better efficient exploration.
Policy Pre-Training via Task-Agnostic Exploration.
More interestingly, we tested the effect of pre-training policies via different objectives as a way to alleviate the well-known hardness of sparse-reward settings, either throught faster learning or zero-short generalization. In order to do so, we employed a multi-agent counterpart of the TRPO algorithm (Schulman et al., 2017) with different pre-trained policies. First, we investigated the effect on the learning curve in the hard-exploration task of Env. (i) under long horizons ($T=150$), with a worst-case goal set on the the opposite corner of the closed room. Pre-training via mixture objectives still lead to a faster learning compared to initializing the policy with a uniform distribution. On the other hand, joint objective pre-training did not lead to substantial improvements over standard initializations. More interestingly, when extremely short horizons were taken into account ($T=50$) the difference became appalling, as shown in Fig. 3(a): pre-training via mixture-based objectives leaded to faster learning and higher performances, while pre-training via disjoint objectives turned out to be even harmful (Fig. 3(b)). This was motivated by the fact that the disjoint objective overfitted the task over the states reachable without coordinated exploration, resulting in almost deterministic policies, as shown in Fig 5 in Appendix B. Finally, we tested the zero-shot capabilities of policy pre-training on the simpler but high dimensional exploration task of Env. (ii), where the goal was sampled randomly between worst-case positions at the boundaries of the region reachable by the arm. As shown in Fig. 4(p), both joint and mixture were able to guarantee zero-shot performances via pre-training compatible with MA-TRPO after learning over $2$e4444 samples, while disjoint objectives were not. On the other hand, pre-training with joint objectives showed an extremely high-variance, leading to worst-case performances not better than the ones of random initialization. Mixture objectives on the other hand showed higher stability in guaranteeing compelling zero-shot performance.
Take-Aways.
Overall, the proposed proof of concepts experiments managed to answer to all of the experimental questions: (a) Algorithm 1 is indeed able to explicitly optimize for finite-trial entropic objectives. Additionally, (b) mixture distributions enforce diverse yet coordinated exploration, that helps when high efficiency is required. Joint or disjoint objectives on the other hand may fail to lead to relevant solutions because of under or over optimization. Finally, (c) efficient exploration enforced by mixture distributions was shown to be a crucial factor not only for the sake of task-agnostic exploration per se, but also for the ability of pre-training via task-agnostic exploration to lead to faster and better training and even zero-shot generalization.
7 Related Works
Below, we summarize the most relevant work investigating multi-agent exploration and task-agnostic exploration in single-agent scenarios.
Multi-Agent Exploration.
Recently, fostering exploration in order to boost performances in (deep) MARL has gained much attention recently. A large set of works proposed to address it via reward-shaping based on many heuristics: Wang et al. (2019) adds a term maximizing the mutual-information between per-agent interactions; Zhang et al. (2021) proposes to optimize the deviation from (jointly) explored regions while Zhang et al. (2023) proposes to optimize directly the entropy over per-agent observations; more recently, Xu et al. (2024) proposed an heuristic reward-shaping enforcing diversity between different agents, and notices that Wang et al. (2019) fails to address the task it introduced. Up to our knowledge, this work is the first in covering both the theoretical properties of multi-agent (task-agnostic) exploration and the optimization of single-trial objectives. Finally, we notice that a that a similar notion of Convex Markov Games was introduced in a concurrent and preliminary work (Gemp et al., 2025), together with some results on existence of equilibria.
Task-Agnostic Exploration and Policy Optimization.
Entropy maximization in MDPs was first introduced in Hazan et al. (2019) and then investigated extensively in a blossoming of different works, addressing the entropy over (trajectories of) states or even observations (to name a few Jin et al., 2020; Golowich et al., 2022; Tiapkin et al., 2023; Zamboni et al., 2024; Savas et al., 2022). Finally, the use of trust-region schemes (Schulman et al., 2017) is ubiquitous in RL. We considered an importance-sampling policy gradient estimator inspired by the work of Metelli et al. (2020). It is yet possible to use other forms of IS estimators, as non-parametric k-NN estimators proposed in Mutti & Restelli (2020).
8 Conclusions and Perspectives
In this paper, we extend the state entropy maximization problem to Markov Games via a novel framework called Convex Markov Games. First of all, we show that the task can be defined in several different ways: one can look at the joint distribution among all the agents, the marginals which are agent-specific, or the mixture which is a tradeoff of the two. Thus, we link these three options via performance bounds and we show that while the first might enjoy nice theoretical guarantees, the others are more promising at working in practice, the latter in particular. Then, we design a practical trust-region algorithm addressing more practical scenarios and we use it to confirm in a set of experiments the expected superiority of mixture objectives, due to its ability to enforce efficient but coordinated exploration over short horizons. Future works can build over our results in many directions, which include pushing forward the known theoretical properties of Convex Markov Games, developing scalable algorithms for continuous domains and investigating more policy classes with succinct representations of the history beyond the one we considered in the experiments. We believe that our work can be a crucial step in the direction of extending state entropy maximization in a principled way to yet more practical settings, in which many agents interact over the same environment.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Appendix A Proofs of the Main Theoretical Results
In this Section, we report the full proofing steps of the Theorems and Lemmas in the main paper.
See 4.1
Proof.
The bounds follow directly from simple yet fundamental relationships between entropies of joint, marginal and mixture distributions which can be found in Paninski (2003); Kolchinsky & Tracey (2017), in particular:
@@ Appendix A Proofs of the Main Theoretical Results
where step (a) and (b) use the fact that $\tilde{d}^{\pi}(s):=\frac{1}{|\mathcal{N}|}\sum_{i\in[\mathcal{N}]}d^{\pi}_{i}(s)$ is a uniform mixture over the agents, whose distribution over the weights has entropy $\log(|\mathcal{N}|)$, so as we can apply the bounds from Kolchinsky & Tracey (2017). Step (c) uses the fact that $H(d^{\pi})=\sum_{i\in[\mathcal{N}]}H(d^{\pi}_{i}|d^{\pi}_{<i})$, then taking the supremum as first $i$ it follows that $\sup_{i\in[\mathcal{N}]}H(d_{i}^{\pi})=H(d^{\pi})-\sum_{j\in[\mathcal{N}]>i}H(d^{\pi}_{j}|d^{\pi}_{<j},d^{\pi}_{i})\leq H(d^{\pi})$ due to non-negativity of entropy. ∎
@@
For the general proof structure, we adapt the steps of Mutti et al. (2023b) for Convex MDPs to the different objectives possible in CMGs. Let us start by considering joint objectives, then:
@@
with $\epsilon>0$ and in step (c) we applied a union bound. We then consider standard concentration inequalities for empirical distributions (Weissman et al., 2003) so to obtain the final bound
@@
A.1 Policy Gradient in CMGs with Infinite-Trials.
In this Section, we analyze policy search for the infinite-trials joint problem $\zeta_{\infty}$ of Eq. (1), via projected gradient ascent over parametrized policies, providing in Th. A.6 the formal counterpart of Fact 4.1 in the Main paper. As a side note, all of the following results hold for the (infinite-trials) mixture objective $\tilde{\zeta}_{\infty}$ of Eq. (5). We will consider the class of parametrized policies with parameters $\theta_{i}\in\Theta_{i}\subset\mathbb{R}^{d}$, with the joint policy then defined as $\pi_{\theta},\theta\in\Theta=\times_{i\in[\mathcal{N}]}\Theta_{i}$. Additionally, we will focus on the computational complexity only, by assuming access to the exact gradient. The study of statistical complexity surpasses the scope of the current work. We define the (independent) Policy Gradient Ascent (PGA) update as:
@@
In general the overall proof is built of three main steps, shared with the theory of Potential Markov Games (Leonardos et al., 2021): (i) prove the existence of well behaved stationary points; (ii) prove that performing independent policy gradient is equivalent to perform joint policy gradient; (iii) prove that the (joint) PGA update converges to the stationary points via single-agent like analysis. In order to derive the subsequent convergence proof, we will make the following assumptions:
Assumption A.1.
@@
On the Validity of Assumption A.1.
This set of assumptions enforces the objective $\zeta_{\infty}(\pi_{\theta})$ to be well-behaved with respect to $\theta$ even if non-convex in general, and will allow for a rather strong result. Yet, the assumptions are known to be true for directly parametrized policies over the whole support of the distribution $d^{\pi}$ (Zhang et al., 2020), and as a result they implicitly require agents to employ policies conditioned over the full state-space $\mathcal{S}$. Fortunately enough, they also guarantee $\Theta$ to be convex.
Lemma A.4 ((i) Global optimality of stationary policies (Zhang et al., 2020)).
Suppose Assumption A.1 holds, and $\mathcal{F}$ is a concave, and continuous function defined in an open neighborhood containing $\lambda(\Theta)$. Let $\theta^{*}$ be a first-order stationary point of problem (1), i.e.,
@@
This result characterizes the optimality of stationary points for Eq. (1). Furthermore, we know from Leonardos et al. (2021) that stationary points of the objective are Nash Equilibria.
Lemma A.5 ((ii) Projection Operator (Leonardos et al., 2021)).
@@
This result will only be used for the sake of the convergence analysis, since it allows to analyze independent updates as joint updates over a single objective. The following Theorem is the formal counterpart of Fact 4.1 and it is a direct adaptation to the multi-agent case of the single-agent proof by Zhang et al. (2020), by exploiting the previous result.
Theorem A.6 ((iii) Convergence rate of independent PGA to stationary points (Formal Fact 4.1)).
Let Assumptions A.1 and A.2 hold. Denote $D_{\lambda}\!:=\!\max_{\lambda,\lambda^{\prime}\in\lambda(\Theta)}\|\lambda-\lambda^{\prime}\|$ as defined in Assumption A.1(iii). Then the independent policy gradient update (8) with $\eta=1/L$ satisfies for all $k$ with respect to a stationary (joint) policy $\pi_{\theta^{*}}$ the following
@@
First, the Lipschitz continuity in Assumption A.2 indicates that
@@
The optimality condition in the policy update rule (8) coupled with the result of Lemma A.5 allows us to follow the same rational as Zhang et al. (2020). We will report their proof structure after this step for completeness.
@@
where step (a) follows from (10) and step (b) uses the convexity of $\lambda(\Theta)$. Then, by the concavity of $\zeta_{\infty}$ and the fact that the composition $\lambda\circ g=id$ due to Assumption A.1(i), we have that:
@@
Moreover, due to Assumption A.1(iii) we have that:
@@
A.2 The Use of Markovian and Non-Markovian Policies in CMGs with Finite-Trials.
The following result describes how in CMGs, as for convex MDPs, Non-Markovian policies are the right policy class to employ to guarantee well-behaved results.
Lemma A.1 (Sufficiency of Disjoint Non-Markvoian Policies).
For every Convex Markov Game $\mathcal{M}$ there exist a joint policy $\pi^{\star}=(\pi^{\star,i})_{i\in\mathcal{N}}$, with $\pi^{\star,i}\in\Delta_{\mathcal{S}^{T}}^{\mathcal{A}^{i}}$ being a deterministic Non-Markovian policy, that is a Nash Equilibrium for non-Disjoint single-trial objectives, for $K=1$.
@@
The proof builds over a straight reduction. We build from the original MG $\mathcal{M}$ a temporally extended Markov Game ~$\tilde{\mathcal{M}}=(\mathcal{N},\tilde{\mathcal{S}},\mathcal{A},\mathbb{P},r,\mu,T)$. A state $\tilde{s}$ is defined for each history that can be induced, i.e., $\tilde{s}\in\tilde{\mathcal{S}}\iff\mathbf{s}\in\mathcal{S}^{T}$. We keep the other objects equivalent, where for the extended transition model we solely consider the last state in the history to define the conditional probability to the next history. We introduce a common reward function across all the agents $r:\tilde{\mathcal{S}}\rightarrow\mathbb{R}$ such that $r(\tilde{s})=H(d(\tilde{s}))$ for joint objectives and $r(\tilde{s})=(1/N)\sum_{i\in[\mathcal{N}]}H(d_{i}(\tilde{s}_{i}))$ for mixture objectives, for all the histories of length T and 00 otherwise. We now know that according to Leonardos et al. (Theorem 3.1, 2021) there exists a deterministic Markovian policy $\tilde{\pi}^{\star}=(\tilde{\pi}^{i})_{i\in\mathcal{N}},\tilde{\pi}^{i}\in\Delta^{\mathcal{A}_{i}}_{\tilde{\mathcal{S}}}$ that is a Nash Equilibrium for $\tilde{\mathcal{M}}$. Since $\tilde{s}$ corresponds to the set of histories of the original game, $\tilde{\pi}^{\star}$ maps to a non-Markovian policy in it. Finally, it is straightforward to notice that the NE of $\tilde{\pi}^{\star}$ for $\tilde{\mathcal{M}}$ implies the NE of $\tilde{\pi}^{\star}$ for the original CMG $\mathcal{M}$. ∎
@@
Lemma A.7 (Behavior of Optimal Markovian Decentralized Policies).
@@
Let us consider the random variable $A_{i}\sim\mathcal{P}_{i}$ denoting the event “the agent $i$ takes action $a^{*}_{i}\in\mathcal{A}_{i}$”. Through the law of total variance (Bertsekas & Tsitsiklis, 2002), we can write the variance of $A$ given $s\in\mathcal{S}$ and $t\geq 0$ as
$\operatorname*{\mathbb{V}\mathrm{ar}}\big{[}A|s,t\big{]}$ $=\operatorname*{\mathbb{E}}\big{[}A^{2}|s,t\big{]}-\operatorname*{\mathbb{E}}\big{[}A|s,t\big{]}^{2}$
$=\operatorname*{\mathbb{E}}_{\mathbf{s}}\Big{[}\operatorname*{\mathbb{E}}\big{[}A^{2}|s,t,\mathbf{s}\big{]}\Big{]}-\operatorname*{\mathbb{E}}_{\mathbf{s}}\Big{[}\operatorname*{\mathbb{E}}\big{[}A|s,t,\mathbf{s}\big{]}\Big{]}^{2}$
@@
We know from Lemma A.1 that the policy $\pi_{\text{NM}}$ is deterministic, so that $\operatorname*{\mathbb{V}\mathrm{ar}}\big{[}\mathcal{B}(\pi_{\text{NM}}(a^{*}|\mathbf{s}\oplus s))\big{]}=0$ for every $\mathbf{s}\oplus s$. We then repeat the same steps in order to compare the two different Markovian policies:
@@
$\operatorname*{\mathbb{V}\mathrm{ar}}\big{[}\mathcal{B}(\pi_{\text{M}}(a^{*}|s_{i},t))\big{]}$ $=\operatorname*{\mathbb{E}}_{\mathbf{s}\oplus(\cdot,s_{-i})\sim p^{\bar{\pi}_{\text{M}}}_{t}}\big{[}\operatorname*{\mathbb{V}\mathrm{ar}}\big{[}\mathcal{B}(\bar{\pi}_{\text{M}}(a^{*}|s_{i},s_{-i},t))\big{]}\big{]}+\operatorname*{\mathbb{V}\mathrm{ar}}_{\mathbf{s}\oplus(\cdot,s_{-i})\sim p^{\bar{\pi}_{\text{M}}}_{t}}\big{[}\operatorname*{\mathbb{E}}\big{[}\mathcal{B}(\bar{\pi}_{\text{M}}(a^{*}|s_{i},s_{-i},t))\big{]}\big{]}$
@@
Appendix B Details on the Experimental Proofs Of Concept.

### v3

@@ Appendix B Proofs of the Main Theoretical Results
where step (a) and (b) use the fact that $\tilde{d}^{\pi}(s):=\frac{1}{|\mathcal{N}|}\sum_{i\in[\mathcal{N}]}d^{\pi}_{i}(s)$ is a uniform mixture over the agents, whose distribution over the weights has entropy $\log(|\mathcal{N}|)$, so as we can apply the bounds from [12]. Step (c) uses the fact that $H(d^{\pi})=\sum_{i\in[\mathcal{N}]}H(d^{\pi}_{i}|d^{\pi}_{<i})$, then taking the supremum as first $i$ it follows that $\sup_{i\in[\mathcal{N}]}H(d_{i}^{\pi})=H(d^{\pi})-\sum_{j\in[\mathcal{N}]>i}H(d^{\pi}_{j}|d^{\pi}_{<j},d^{\pi}_{i})\leq H(d^{\pi})$ due to non-negativity of entropy. ∎
@@
For the general proof structure, we adapt the steps of [26] for cMDPs to the different objectives possible in cMGs. Let us start by considering joint objectives, then:
@@
with $\epsilon>0$ and in step (c) we applied a union bound. We then consider standard concentration inequalities for empirical distributions [40] so to obtain the final bound
@@
B.1 Policy Gradient in cMGs with Infinite-Trials Formulations.
In this Section, we analyze policy search for the infinite-trials joint problem $\zeta_{\infty}$ of Eq. (1), via projected gradient ascent over parametrized policies, providing in Th. B.6 the formal counterpart of Fact 4.1 in the Main paper. As a side note, all of the following results hold for the (infinite-trials) mixture objective $\tilde{\zeta}_{\infty}$ of Eq. (3). We will consider the class of parametrized policies with parameters $\theta_{i}\in\Theta_{i}\subset\mathbb{R}^{d}$, with the joint policy then defined as $\pi_{\theta},\theta\in\Theta=\times_{i\in[\mathcal{N}]}\Theta_{i}$. Additionally, we will focus on the computational complexity only, by assuming access to the exact gradient. The study of statistical complexity surpasses the scope of the current work. We define the (independent) Policy Gradient Ascent (PGA) update as:
@@
In general the overall proof is built of three main steps, shared with the theory of Potential Markov Games [15]: (i) prove the existence of well behaved stationary points; (ii) prove that performing independent policy gradient is equivalent to perform joint policy gradient; (iii) prove that the (joint) PGA update converges to the stationary points via single-agent like analysis. In order to derive the subsequent convergence proof, we will make the following assumptions:
Assumption B.1.
@@
On the Validity of Assumption B.1.
This set of assumptions enforces the objective $\zeta_{\infty}(\pi_{\theta})$ to be well-behaved with respect to $\theta$ even if non-convex in general, and will allow for a rather strong result. Yet, the assumptions are known to be true for directly parametrized policies over the whole support of the distribution $d^{\pi}$ [48], and as a result they implicitly require agents to employ policies conditioned over the full state-space $\mathcal{S}$. Fortunately enough, they also guarantee $\Theta$ to be convex.
Lemma B.4 ((i) Global optimality of stationary policies [48]).
Suppose Assumption B.1 holds, and $\mathcal{F}$ is a concave, and continuous function defined in an open neighborhood containing $\lambda(\Theta)$. Let $\theta^{*}$ be a first-order stationary point of problem (1), i.e.,
@@
This result characterizes the optimality of stationary points for Eq. (1). Furthermore, we know from [15] that stationary points of the objective are Nash Equilibria.
Lemma B.5 ((ii) Projection Operator [15]).
@@
This result will only be used for the sake of the convergence analysis, since it allows to analyze independent updates as joint updates over a single objective. The following Theorem is the formal counterpart of Fact 4.1 and it is a direct adaptation to the multi-agent case of the single-agent proof by [48], by exploiting the previous result.
Theorem B.6 ((iii) Convergence rate of independent PGA to stationary points (Formal Fact 4.1)).
Let Assumptions B.1 and B.2 hold. Denote $D_{\lambda}\!:=\!\max_{\lambda,\lambda^{\prime}\in\lambda(\Theta)}\|\lambda-\lambda^{\prime}\|$ as defined in Assumption B.1(iii). Then the independent policy gradient update (5) with $\eta=1/L$ satisfies for all $k$ with respect to a stationary (joint) policy $\pi_{\theta^{*}}$ the following
@@
First, the Lipschitz continuity in Assumption B.2 indicates that
@@
The optimality condition in the policy update rule (5) coupled with the result of Lemma B.5 allows us to follow the same rational as [48]. We will report their proof structure after this step for completeness.
@@
where step (a) follows from (7) and step (b) uses the convexity of $\lambda(\Theta)$. Then, by the concavity of $\zeta_{\infty}$ and the fact that the composition $\lambda\circ g=id$ due to Assumption B.1(i), we have that:
@@
Moreover, due to Assumption B.1(iii) we have that:
@@
B.2 The Use of Markovian and Non-Markovian Policies in cMGs with Finite-Trials Formulations.
The following result describes how in cMGs, as for cMDPs, Non-Markovian policies are the right policy class to employ to guarantee well-behaved results.
Lemma B.1 (Sufficiency of Disjoint Non-Markvoian Policies).
For every cMG $\mathcal{M}$ there exist a joint policy $\pi^{\star}=(\pi^{\star,i})_{i\in\mathcal{N}}$, with $\pi^{\star,i}\in\Delta_{\mathcal{S}^{T}}^{\mathcal{A}^{i}}$ being a deterministic Non-Markovian policy, that is a Nash Equilibrium for non-Disjoint single-trial objectives, for $K=1$.
@@
The proof builds over a straight reduction. We build from the original MG $\mathcal{M}$ a temporally extended Markov Game ~$\tilde{\mathcal{M}}=(\mathcal{N},\tilde{\mathcal{S}},\mathcal{A},\mathbb{P},r,\mu,T)$. A state $\tilde{s}$ is defined for each history that can be induced, i.e., $\tilde{s}\in\tilde{\mathcal{S}}\iff\mathbf{s}\in\mathcal{S}^{T}$. We keep the other objects equivalent, where for the extended transition model we solely consider the last state in the history to define the conditional probability to the next history. We introduce a common reward function across all the agents $r:\tilde{\mathcal{S}}\rightarrow\mathbb{R}$ such that $r(\tilde{s})=H(d(\tilde{s}))$ for joint objectives and $r(\tilde{s})=(1/N)\sum_{i\in[\mathcal{N}]}H(d_{i}(\tilde{s}_{i}))$ for mixture objectives, for all the histories of length T and 00 otherwise. We now know that according to [Theorem 3.1, 15] there exists a deterministic Markovian policy $\tilde{\pi}^{\star}=(\tilde{\pi}^{i})_{i\in\mathcal{N}},\tilde{\pi}^{i}\in\Delta^{\mathcal{A}_{i}}_{\tilde{\mathcal{S}}}$ that is a Nash Equilibrium for $\tilde{\mathcal{M}}$. Since $\tilde{s}$ corresponds to the set of histories of the original game, $\tilde{\pi}^{\star}$ maps to a non-Markovian policy in it. Finally, it is straightforward to notice that the NE of $\tilde{\pi}^{\star}$ for $\tilde{\mathcal{M}}$ implies the NE of $\tilde{\pi}^{\star}$ for the original cMG $\mathcal{M}$. ∎
@@
Lemma B.7 (Behavior of Optimal Markovian Decentralized Policies).
@@
Let us consider the random variable $A_{i}\sim\mathcal{P}_{i}$ denoting the event “the agent $i$ takes action $a^{*}_{i}\in\mathcal{A}_{i}$”. Through the law of total variance [3], we can write the variance of $A$ given $s\in\mathcal{S}$ and $t\geq 0$ as
$\operatorname*{\mathbb{V}\mathrm{ar}}\big{[}A|s,t\big{]}$ $=\operatorname*{\mathbb{E}}\big{[}A^{2}|s,t\big{]}-\operatorname*{\mathbb{E}}\big{[}A|s,t\big{]}^{2}=\operatorname*{\mathbb{E}}_{\mathbf{s}}\Big{[}\operatorname*{\mathbb{E}}\big{[}A^{2}|s,t,\mathbf{s}\big{]}\Big{]}-\operatorname*{\mathbb{E}}_{\mathbf{s}}\Big{[}\operatorname*{\mathbb{E}}\big{[}A|s,t,\mathbf{s}\big{]}\Big{]}^{2}$
@@
We know from Lemma B.1 that the policy $\pi_{\text{NM}}$ is deterministic, so that $\operatorname*{\mathbb{V}\mathrm{ar}}\big{[}\mathcal{B}(\pi_{\text{NM}}(a^{*}|\mathbf{s}\oplus s))\big{]}=0$ for every $\mathbf{s}\oplus s$. We then repeat the same steps in order to compare the two different Markovian policies:
@@
$\operatorname*{\mathbb{V}\mathrm{ar}}\big{[}\mathcal{B}(\pi_{\text{M}}(a^{*}|s_{i},t))\big{]}$ $=\operatorname*{\mathbb{E}}_{\mathbf{s}\oplus(\cdot,s_{-i})\sim p^{\bar{\pi}_{\text{M}}}_{t}}\Big{[}\operatorname*{\mathbb{V}\mathrm{ar}}\big{[}\mathcal{B}(\bar{\pi}_{\text{M}}(a^{*}|s_{i},s_{-i},t))\big{]}+\operatorname*{\mathbb{E}}\big{[}\mathcal{B}(\bar{\pi}_{\text{M}}(a^{*}|s_{i},s_{-i},t))\big{]}\Big{]}$
@@
Appendix C Details on the Empirical Corroboration.
All the experiments were performed over an Apple M2 chip (8-core CPU, 8-core GPU, 16-core Neural Engine) with 8 GB unified memory with a maximum time of execution of 24 hours.